Quantitative Problem: Barton Industries estimates its cost of common equity by using three approaches: the CAPM, the bond-yield-plus-risk-premium approach, and the DCF model. Barton expects next year's annual dividend, D1, to be $1.90 and it expects dividends to grow at a constant rate g = 3.6%. The firm's current common stock price, P0, is $25.00. The current risk-free rate, rRF, = 4.9%; the market risk premium, RPM, = 6.5%, and the firm's stock has a current beta, b, = 1.35. Assume that the firm's cost of debt, rd, is 5.22%. The firm uses a 3.5% risk premium when arriving at a ballpark estimate of its cost of equity using the bond-yield-plus-risk-premium approach. What is the firm's cost of equity using each of these three approaches? Round your answers to two decimal places.

CAPM cost of equity:   %
Bond yield plus risk premium:   %
DCF cost of equity:   %

What is your best estimate of the firm's cost of equity?
Answer #1

CAPM cost of equity= risk free rate+beta*RPm=4.9%+1.35*6.5%=13.68%

Bond yield plus risk premium=5.22%+3.5%=8.72%

DCF cost of equity=(1.90/25)+3.6%=11.20%

best estimate=(13.68%+8.72%+11.20%)/3=11.20%